From 1990 to 2000, the population of Texas increased from 16,985 to 20,851,820. If the growth rate remained same, what would be the population of Texas in 2030?
Answer by checkley71(8403)

(20,851,820-16,985)/10=20,834,835/10=20,834,84 INCREASE PER YEAR BETWEEN 1990 & 2000.
AT THIS RATE THE NEXT 30 YEARS WILL BE:
20,851,820+30*20,834,835=20,851,820+625,045,050=645,896,870 ANSWER.